OpenStudy (anonymous):

Can you simplify an exponent within a logartithm? Ex Ln x^7 - Ln x^ 5, does it equal Ln x ^2 or something else's

OpenStudy (loser66):

yes = ln x^2

OpenStudy (anonymous):

Doesn't it affexct the whole problem though

OpenStudy (loser66):

no, it's the way you calculate log. log a + log b = log (a*b) and log a - log b = log (a/b) .

OpenStudy (anonymous):

Do you know how to calculate magnitude with. Given angle